If g(x) = 3x - 8, find g[g(-4)].
Thanks

Found 2 solutions by fractalier, Nate:
Answer by fractalier(6550)

g(g(-4)) = g(3(-4) - 8) = g(-20) = 3(-20) - 8 = -68

Answer by Nate(3500)

g(x) = 3x - 8
[g o g] = 3(3x - 8) - 8 = 9x - 32
g[g(-4)] = 9x - 32 = 9(-4) - 32 = -36 - 32 = -68